I have a garden that is 20m by 12m and a path around the garden that is 2m wide. What is the area of the path?

Answer by checkley79(3341)

AREA=L*W
AREA=2(20+2)+2(12+2)
AREA=2*22+2*14
AREA=44+28
AREA=72 M^2.